OpenStudy (anonymous):

sinx/1-cotx=cosx/1-tanx=cosx+sinx

OpenStudy (mathmath333):

is this the question \(\large\tt \begin{align} \color{black}{\dfrac{sinx}{1-cotx}=\dfrac{cosx}{1-tanx}=cosx+sinx }\end{align}\)